OpenStudy (starryknight):

Solve for y in the following system of equations. -8x - 8y = -64 8x - 8y = -96 -10 16 10 8 please help me whoever gives me the best answer gets a free metal

OpenStudy (anonymous):

y = 10

OpenStudy (mwhit2101):

-8x-9y=-64---Y=8-x 8x-8y=-96----y=x+12